# From Coarse to Fine-Grained Emotion Annotation: An Immediate Recall Paradigm with Validation through Physiological Evidence and Recognition Performance

### v1

@@ III-B Participants
A total of twenty healthy participants (12 males, 8 females) were recruited for this study. Their ages ranged from 22 to 30 years, with a mean age of 24.7 years ± 2.1 years (Mean ± SD). All participants reported being right-handed, having normal or corrected-to-normal vision, and reported no history of neurological or psychiatric disorders. This study received ethical approval from the Northwestern Polytechnical University Ethics Committee. Before the experiment, the procedure and purpose of the study were introduced to the participants. Subsequently, by signing a consent agreement, participants agreed to the recording, sharing, and publication of their physiological signals and facial images for academic purposes. Participants were aware that their participation was voluntary and that they were free to withdraw at any time without providing any reason.
@@ III-C Protocol
(3) Immediate Replay and Annotation: Immediately following stimulus presentation, the video was replayed. Participants recalled their emotional experience from the initial viewing and indicated to the experimenter when they first experienced the target emotion. Due to GSR and PPG sensors on their fingers preventing direct interface control, the experimenter operated the annotation software based on participant instructions. Using the interface controls shown in Fig. 3, including an adjustable progress bar, the experimenter navigated to the precise timestamp identified by the participant. Upon participant confirmation, the experimenter paused the replay and recorded the timestamp. Participants then reported the emotion label and intensity level (Low, Medium, High), which the experimenter entered into the annotation panel, thereby linking the temporal, categorical, and intensity dimensions of the emotional response.
@@ III-D Data preprocessing
Prior to analysis, the continuously recorded raw multimodal signals underwent filtering to remove noise and artifacts by MNE-Python [38]. Specifically:
EEG data: The EEG signals were first downsampled to 250 Hz to reduce computational load while preserving relevant frequency information. The data were then band-pass filtered between 0.5 Hz and 70 Hz using a finite impulse response (FIR) filter, followed by a 50 Hz notch filter to remove power line noise. Independent Component Analysis (ICA) [39] was employed to identify and remove typical artifacts including eye blinks (EOG) and muscle activity (EMG). Bad channels were identified using a variance-based criterion: channels with variance exceeding 10 times the median variance across all channels were marked as ’bad’ and subsequently repaired using spherical spline interpolation.
@@
Subsequently, analysis epochs were extracted based on the discrete event timestamps $t_{event}$ identified during the annotation phase. Acknowledging that $t_{event}$ represents a subjective recollection of emotion onset and thus likely contains temporal jitter (due to factors such as memory precision limitations, reaction time during annotation, and the inherent difficulty in pinpointing the exact onset of a subjective feeling), and considering that physiological responses typically unfold over a period rather than being instantaneous, we employed a fixed window approach for analysis. Specifically, for each annotated $t_{event}$, a 4-second window was extracted from the preprocessed multimodal data streams [40], encompassing 2 seconds before and 2 seconds after $t_{event}$ (i.e., [$t_{event}-2\,\text{s},t_{event}+2\,\text{s}$]). This window duration was selected to capture the most salient physiological changes surrounding the reported emotional event while accommodating the potential temporal variability [41].
@@ IV Physiological Validation of the Annotation Paradigm
In Section III, we detailed the experimental procedure, the novel annotation paradigm, and the preprocessing methods for the multimodal data. This Section presents in-depth analyses of the multimodal physiological signals based on the processed data, the corresponding annotations, and the time windows defined around the annotated events, aiming to physiologically validate the annotation effectiveness. The analysis comprises two main parts: first, the results of the EEG signal analysis in Section IV-A; second, the results of the peripheral physiological signal analysis in Section IV-B; and finally, Section IV-C quantitatively evaluates the comparative results of our immediate annotation paradigm relative to a delayed annotation paradigm.
@@ IV-A EEG Analysis Results
This subsection details the EEG analysis results aimed at identifying brain neural correlates associated with the precisely annotated emotional event markers $t_{event}$ derived from the FIRMED. We focused on event-related changes in Power Spectral Density (PSD), comparing activity within the emotion-event time windows to the pre-trial baseline period [42]. PSD was chosen as it is a well-established method for quantifying the distribution of power across different frequency bands in EEG signals, providing insights into the neural oscillatory changes that are widely recognized as correlates of various cognitive and emotional states. We calculated the PSD for each channel across standard frequency bands: $\delta$ (1-4 Hz), $\theta$ (4-8 Hz), $\alpha$ (8-13 Hz), $\beta$ (13-30 Hz) and $\gamma$ (30-45 Hz). For each participant and trial, preprocessed EEG data were extracted for the emotion-event time windows and the 10s baseline window. PSD for each channel within each window was calculated using Welch’s method.
For each participant and emotion category, the difference change in PSD from baseline to the event window is shown in (1). Subsequently, to assess the statistical significance of these changes at the group level, one-sample t-tests were performed at each electrode location to test whether the group-averaged $\Delta PSD$ values differed significantly from zero. To address the multiple comparisons problem inherent in simultaneous testing across 59 EEG channels, we employed cluster-based permutation testing [43]. This approach involves forming spatial clusters of electrodes showing consistent effects based on their adjacency relationships, followed by 1000 random permutations where the sign of each participant’s $\Delta PSD$ values is randomly flipped to simulate the null hypothesis of no systematic effect. Cluster-level statistics (sum of t-values within each cluster) from the observed data are then compared against the distribution of maximum cluster statistics obtained from the permutations to determine statistical significance. This non-parametric statistical method effectively controls the family-wise error rate by taking into account the spatial dependencies between channels, thereby offering greater sensitivity to detect genuine, spatially coherent neural effects. The findings described below highlight patterns observed at the group level based on these average effects.
@@
Happiness: Group-level analysis of happiness events revealed widespread neural oscillatory changes characterized by multi-frequency engagement as shown in Fig. 4. The $\alpha$ band demonstrated the most extensive response, with 48 of 59 electrodes (81.4%) forming a single significant cluster ($p=0.004$, $d=-0.63$ to $0.03$). This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of positive emotional states [44]. The $\gamma$ band showed remarkable synchronization, with 44 electrodes (74.6%) displaying significant power increases within a coherent cluster ($p=0.001$, $d=0.00$ to $1.11$). This extensive $\gamma$ enhancement, particularly prominent in frontal, temporal, parietal, and occipital regions, supports the neural binding hypothesis and suggests widespread cross-regional integration during positive emotional experiences [45]. The $\beta$ band exhibited a selective activation pattern involving 27 electrodes (45.8%) organized into two distinct clusters ($p=0.018$ and $p=0.001$, $d=-0.46$ to $1.17$). This bi-cluster organization likely reflects coordinated activation of prefrontal-parietal cognitive control and temporo-occipital emotion regulation networks.
Figure 4: Group-level statistical T-maps for the happiness events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
In contrast, lower frequency bands showed markedly different response patterns. The $\delta$ frequency band demonstrated limited but focused significant changes affecting only 5 out of 59 channels, primarily localized to left temporal and parieto-occipital region. The $\theta$ frequency band did not exhibit any statistically significant clustered effects, indicating that this frequency range did not show spatially coherent changes during happy emotion processing in this study.
Surprise: Surprise events demonstrated more constrained cortical activation compared to other emotions, with effects primarily confined to mid-to-high frequency bands as shown in Fig. 5. The $\alpha$ band showed the most prominent response, with 32 of 59 electrodes (54.2%) forming two distinct clusters ($p=0.011$ and $p=0.001$, $d=-0.87$ to $0.04$). This $\alpha$ desynchronization, primarily involving central sensorimotor and posterior parietal regions, was substantially more limited than the near-global patterns observed for emotions like happinsess, suggesting a more focused cortical response. The $\beta$ band exhibited selective activation with only 11 electrodes (18.6%) organized into two significant clusters ($p=0.015$ and $p=0.001$, $d=-0.65$ to $0.23$), likely reflecting limited engagement of frontal and parietal cognitive control regions necessary for rapid reorientation. The $\gamma$ band showed minimal response, with merely 5 electrodes (8.5%) forming a single cluster ($p=0.018$, $d=-0.53$ to $0.29$), indicating restricted high-frequency synchronization in specific attention-related areas.
Notably, both $\delta$ and $\theta$ bands showed no statistically significant clustered effects, clearly distinguishing surprise from emotions that engage low-frequency oscillations. This absence of low-frequency involvement, combined with moderate $\alpha$ desynchronization and limited high-frequency engagement, suggests that surprise represents a relatively simple and transient emotional state primarily characterized by rapid attentional reorientation rather than the deep neural integration observed in more complex emotional experiences [46].
Figure 5: Group-level statistical T-maps for the surprise events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Disgust: Group-level analysis of Disgust events evealed a distinctive neurophysiological profile characterized by selective frequency-band modulations, most notably featuring $\theta$ band suppression as shown in Fig. 6. A unique finding was the significant power decrease in the $\theta$ band, localized to a cluster of 5 electrodes (8.5%) in the right posterior temporal-parietal-occipital region ($p=0.044$, $d=-1.04$ to $0.61$). This localized $\theta$ suppression, absent in all other analyzed emotions, may reflect specific inhibitory processes and attentional disengagement mechanisms associated with aversive stimulus processing. The $\alpha$ band demonstrated significant and widespread desynchronization, with 27 electrodes (45.8%) forming a single cluster ($p=0.025$). Notably, effect sizes were uniformly negative ($d=-0.77$ to $-0.07$), indicating consistent cortical activation across affected regions, primarily in central and posterior areas. This pattern suggests moderate but reliable cortical engagement during disgust processing, though less extensive than observed for high-arousal emotions. The $\beta$ band showed selective modulation involving 14 electrodes (23.7%) in a significant cluster ($p=0.003$, $d=-0.94$ to $0.70$), concentrated in parietal and parieto-occipital regions. The mixed positive and negative effect sizes suggest heterogeneous cortical responses, potentially reflecting the complex interplay between sensory rejection and motor preparation for avoidance behaviors.
No statistically significant clusters were identified in the $\delta$ or $\gamma$ bands after correction for multiple comparisons. This absence of low-frequency slow-wave activity and high-frequency binding distinguishes Disgust from emotions like Fear (characterized by $\delta$ enhancement) and Happy (marked by $\gamma$ synchronization). The distinct combination of $\theta$ suppression, moderate $\alpha$ desynchronization, and selective $\beta$ modulation constitutes the unique neurophysiological signature of disgust.
Figure 6: Group-level statistical T-maps for the disgust events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Anger: Group-level analysis of anger events revealed the most restricted cortical activation pattern among all analyzed emotions, with significant modulations primarily confined to the $\alpha$ band as shown in Fig. 7. The $\alpha$ band showed moderate desynchronization, with 21 of 59 electrodes (35.6%) forming a single significant cluster ($p=0.004$, $d=-0.87$ to $-0.002$), concentrated over central, parietal, and occipital regions. This $\alpha$ suppression, while statistically significant, was markedly less extensive compared to emotions like happiness, suggesting a more localized cortical arousal pattern consistent with the focused nature of anger-related processing. The $\delta$ band demonstrated very focal activation with only 2 electrodes (3.4%) forming a significant cluster ($p=0.042$, $d=-0.52$ to $0.23$) in the left frontal area. Similarly, the $\beta$ band showed highly restricted activation involving merely 3 electrodes (5.1%) in a significant cluster ($p=0.035$, $d=-0.80$ to $0.24$), localized to right temporal regions.
Notably, both $\theta$ and $\gamma$ bands showed no statistically significant clustered effects after correction for multiple comparisons. The neural signature of anger in this paradigm was characterized by moderate $\alpha$ desynchronization with minimal involvement of other frequency bands, suggesting that anger may have a more restricted or individual-specific cortical response pattern compared to other basic emotions.
Figure 7: Group-level statistical T-maps for the anger events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Fear: Group-level analysis of fear events revealed profound cortical modulations characterized by multi-frequency engagement as shown in Fig. 8. The $\alpha$ band demonstrated the most extensive response, with 57 electrodes (96.6%) forming a single significant cluster ($p=0.001$, $d=-1.105$ to $-0.136$). This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of threat detection states. Remarkably, the $\delta$ band showed substantial power increases, with 25 electrodes (42.4%) distributed across two significant clusters ($p=0.010$ and $p=0.026$, $d=0.222$ to $0.638$), primarily in frontal and temporal regions. This extensive delta synchronization, unique among all analyzed emotions, suggests deep engagement of limbic-cortical circuits involved in threat processing and emotional memory consolidation. The $\beta$ band exhibited significant modulation with 20 electrodes (33.9%) forming a single cluster ($p=0.001$, $d=-1.370$ to $0.288$), likely reflecting enhanced cognitive control and motor preparation networks.
In contrast, neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons, distinguishing fear from emotions that engage these higher-frequency oscillations threat response.
Figure 8: Group-level statistical T-maps for the fear events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Sadness: Group-level analysis of sadness events revealed widespread cortical activation characterized by prominent mid-frequency modulations as shown in Fig. 9. The $\alpha$ band demonstrated extensive desynchronization with 45 out of 59 electrodes (76.3%) forming a single significant cluster ($p=0.004$, $d=-0.772$ to $-0.258$). This widespread $\alpha$ suppression, encompassing frontal, central, temporal, parietal, and occipital regions, indicates comprehensive cortical disinhibition and heightened neural processing associated with sadness. The $\beta$ band showed moderate but significant activation with 12 electrodes (20.3%) forming a cluster ($p=0.005$, $d=-0.623$ to $0.173$), concentrated in parietal and parieto-occipital regions, likely reflecting cognitive-emotional regulation processes. Notably, the $\delta$ band showed highly focal activation with only 1 electrode (1.7%) at F1 reaching significance ($p=0.015$, $d=-0.505$ to $-0.169$), suggesting limited involvement of deep emotional processing circuits compared to high-arousal emotions like fear.
Neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons. The predominant $\alpha$ desynchronization coupled with selective $\beta$ engagement and minimal low-frequency involvement constitutes the characteristic neurophysiological signature of sadness.
Figure 9: Group-level statistical T-maps for the sadness events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
In summary, our research systematically delineates the unique neurophysiological signatures of six basic emotions. A common pattern observed was widespread $\alpha$ band desynchronization, which, acting as a general index of cortical arousal, was present across all emotions. Building upon this, several emotions exhibited characteristic spectral patterns dominated by specific frequency bands: Fear was uniquely defined by a concurrent, profound $\alpha$ band suppression and widespread $\delta$ band synchronization; Happiness was distinguished from other emotions by its hallmark large-scale $\gamma$ band synchronization; while Disgust was identified by its unique marker of localized $\theta$ band suppression in the right posterior region. Meanwhile, other emotions displayed more subtle or mixed patterns: Sadness was characterized by broad desynchronization across both the $\alpha$ and $\beta$ bands; the response elicited by Surprise was more topographically constrained; and Anger produced the weakest and most localized PSD changes, indicating its cortical signature is less distinct under the current paradigm. The ability of our study to consistently resolve these multi-level neural oscillatory patterns—which are also corroborated by existing neuroscience research—ultimately confirms the effectiveness of the FIRMED precise annotation paradigm in capturing and differentiating genuine, fine-grained neurophysiological responses of emotion.

### v2

@@ III-B Participants
A total of thirty-five healthy participants (20 males, 15 females) were recruited for this study. Their ages ranged from 22 to 30 years, with a mean age of 23.9 years ± 1.7 years (Mean ± SD). All participants reported being right-handed, having normal or corrected-to-normal vision, and reported no history of neurological or psychiatric disorders. This study received ethical approval from the Northwestern Polytechnical University Ethics Committee. Before the experiment, the procedure and purpose of the study were introduced to the participants. Subsequently, by signing a consent agreement, participants agreed to the recording, sharing, and publication of their physiological signals and facial images for academic purposes. Participants were aware that their participation was voluntary and that they were free to withdraw at any time without providing any reason.
@@ III-C Protocol
(3) Immediate Replay and Annotation: Immediately following stimulus presentation, the video was replayed. Participants were instructed to recall their emotional experience from the initial viewing and indicate to the experimenter each salient moment they experienced the target emotion during the replay. Due to GSR and PPG sensors on their fingers preventing direct interface control, the experimenter operated the annotation software in Fig. 3 based on participant instructions. Using interface controls with adjustable progress bar, the experimenter navigated to the precise timestamp identified by the participant for each reported emotional event. Upon participant confirmation for a specific event, the experimenter paused the replay and recorded the timestamp ($t_{event}$). The participant then reported the corresponding emotion label and intensity level (Low, Medium, High) for that specific event, which the experimenter entered into the annotation panel. This process allowed for the marking of multiple timestamps, labels, and intensities within a single trial.
@@ III-D Data preprocessing
Prior to analysis, the continuously recorded raw multimodal signals underwent filtering to remove noise and artifacts by MNE-Python [gramfort2014mne]. Specifically:
EEG data: The EEG signals were first downsampled to 250 Hz to reduce computational load while preserving relevant frequency information. The data were then band-pass filtered between 0.5 Hz and 70 Hz using a finite impulse response (FIR) filter, followed by a 50 Hz notch filter to remove power line noise. Independent Component Analysis (ICA) [makeig1995independent] was employed to identify and remove typical artifacts including eye blinks (EOG) and muscle activity (EMG). Bad channels were identified using a variance-based criterion: channels with variance exceeding 10 times the median variance across all channels were marked as ’bad’ and subsequently repaired using spherical spline interpolation.
@@
Subsequently, analysis epochs were extracted based on the discrete event timestamps $t_{event}$ identified during the annotation phase. This $t_{event}$, captured via immediate recall, represents the participant’s most direct and fresh recollection of the peak emotional moment, offering a vastly more precise temporal anchor than traditional whole-trial annotation methods. We then employed a fixed 4-second window centered around this high-fidelity timestamp ($t_{event}-2\,\text{s},t_{event}+2\,\text{s}$]) for analysis [hajcak2010event]. This approach is a standard and pragmatic choice in event-related physiological analysis; while acknowledging the potential for minor temporal jitter in subjective reports, this window is designed to capture the salient physiological activity occurring around the reported peak, thereby inherently accommodating such jitter and the natural unfolding of the physiological response [muhl2014survey].
@@ IV Physiological Validation of the Annotation Paradigm
In Section III, we detailed the experimental procedure, our novel annotation paradigm, and the data preprocessing methods. This Section now presents the in-depth physiological analyses designed to validate the effectiveness of our annotation method by establishing a robust correspondence between subjective reports and objective physiological events.
To first provide a holistic and intuitive demonstration of this correspondence, we present a representative physiological recording from a high-arousal ”Fear” trial in Fig. 4. As is visually evident, these subjectively annotated timestamps align perfectly with dramatic, multi-system physiological responses. The EEG exhibits large, synchronous slow-wave activity, while the GSR signal initiates a steep ascent after the first event and reaches its absolute peak after the second.
Figure 4: Screenshots of real-time recordings of multisystem physiological responses during a fear experiment, showing baseline, unlabeled data, and labeled ”fear” data. EEG, GSR, and ECG show minimal fluctuations during the baseline and unlabeled periods. In contrast, significant and synchronized physiological responses are specifically locked to the subjectively labeled moments of fear ($t_{\text{event}_{1}}$, $t_{\text{event}_{2}}$), demonstrating the temporal precision of the annotation method.
This striking alignment between subjective annotation and the objective, multi-system physiological narrative serves as the primary motivation for the detailed analyses that follow. In the subsequent subsections, we will systematically and quantitatively dissect these phenomena, first examining the neural correlates in the CNS in Section IV-A, then the arousal-based responses in the ANS in Section IV-B, and finally, through a direct comparison with a delayed recall paradigm in Section IV-C.
@@ IV-A EEG Analysis Results
This subsection details the EEG analysis results aimed at identifying brain neural correlates associated with the precisely annotated emotional event markers $t_{event}$ derived from the FIRMED. We focused on event-related changes in Power Spectral Density (PSD), comparing activity within the emotion-event time windows to the baseline period [liu2017real]. PSD was chosen as it is a well-established method for quantifying the distribution of power across different frequency bands in EEG signals, providing insights into the neural oscillatory changes that are widely recognized as correlates of various cognitive and emotional states. We calculated the PSD for each channel across standard frequency bands: $\delta$ (1-4 Hz), $\theta$ (4-8 Hz), $\alpha$ (8-13 Hz), $\beta$ (13-30 Hz) and $\gamma$ (30-45 Hz). For each participant and trial, preprocessed EEG data were extracted for the emotion-event time windows and the 10s baseline window. PSD for each channel within each window was calculated using Welch’s method.
For each participant and emotion category, the difference change in PSD from baseline to the event window is shown in (1). Subsequently, to assess the statistical significance of these changes at the group level, one-sample t-tests were performed at each electrode location to test whether the group-averaged $\Delta PSD$ values differed significantly from zero. To address the multiple comparisons problem inherent in simultaneous testing across 59 EEG channels, we employed cluster-based permutation testing [maris2007nonparametric]. This approach involves forming spatial clusters of electrodes showing consistent effects based on their adjacency relationships, followed by 1000 random permutations where the sign of each participant’s $\Delta PSD$ values is randomly flipped to simulate the null hypothesis of no systematic effect. Cluster-level statistics (sum of t-values within each cluster) from the observed data are then compared against the distribution of maximum cluster statistics obtained from the permutations to determine statistical significance. This non-parametric statistical method effectively controls the family-wise error rate by taking into account the spatial dependencies between channels, thereby offering greater sensitivity to detect genuine, spatially coherent neural effects. The findings described below highlight patterns observed at the group level based on these average effects can be seen in Fig. 5 and Table II.
@@
Figure 5: Event-related EEG PSD change T-maps (4s window) for six emotions across five frequency bands. White circles indicate channels in significant clusters (cluster permutation test, $p<0.05$). Red/blue indicates PSD increase/decrease. The figure compares neural signatures across emotions.
TABLE II: EEG Neurophysiological Signatures of Six Basic Emotions: Frequency-Specific Power Modulation and Regional Distribution Patterns
Emotion Key Frequency Bands Key Brain Regions
Happiness $\alpha$ (8–13 Hz): $\downarrow$ 67.8% $\gamma$ (30–45 Hz): $\uparrow$ 28.8% $\beta$ (13–30 Hz): $\uparrow$ 10.2% $\bullet$ Frontal-parietal: $\alpha$ desynchronization $\bullet$ Temporal-frontal: $\gamma$, $\beta$ synchronization
Surprise $\alpha$ (8–13 Hz): $\downarrow$ 35.6% $\beta$ (13–30 Hz): $\downarrow$ 35.6% $\gamma$ (30–45 Hz): $\downarrow$ 27.1% $\bullet$ Parietal-occipital: $\alpha$ desynchronization $\bullet$ Temporal-parietal: $\beta$, $\gamma$ desynchronization
Disgust $\alpha$ (8–13 Hz): $\downarrow$ 25.4% $\theta$ (4–8 Hz): $\downarrow$ 10.2% $\bullet$ Frontal-central-parietal: $\alpha$ desynchronization $\bullet$ Frontal: $\theta$ desynchronization
Anger $\alpha$ (8–13 Hz): $\downarrow$ 45.8% $\beta$ (13–30 Hz): $\downarrow$ 15.3% $\bullet$ Frontal-temporal-central-parietal: $\alpha$ desynchronization $\bullet$ Temporal-central-parietal: $\beta$ desynchronization
Fear $\alpha$ (8–13 Hz): $\downarrow$ 91.5% $\beta$ (13–30 Hz): $\downarrow$ 30.5% $\theta$ (4–8 Hz): $\downarrow$ 20.3% $\bullet$ Frontal-temporal-central-parietal: $\alpha$ desynchronization $\bullet$ Frontal-central-parietal: $\beta$ desynchronization $\bullet$ Central-parietal: $\theta$ desynchronization
Sadness $\alpha$ (8–13 Hz): $\downarrow$ 81.4% $\beta$ (13–30 Hz): $\downarrow$ 39.0% $\gamma$ (30–45 Hz): $\downarrow$ 16.9% $\bullet$ Frontal-temporal-parietal-occipital: $\alpha$ desynchronization $\bullet$ Frontal-temporal-parietal: $\beta$ desynchronization $\bullet$ Parietal-occipital: $\gamma$ desynchronization
[…]
Happiness: Happiness emotion elicited moderate but significant $\alpha$ desynchronization across 40 of electrodes ($p<0.001$, d=0.56d=-0.56 to 0.29-0.29), distributed in frontal and parietal regions, indicating enhanced cortical engagement. Significant $\gamma$ band enhancement was observed across 17 of electrodes ($p<0.05$, $d=0.30$ to $0.63$), primarily in temporal and frontal areas, reflecting heightened high-frequency neural synchronization associated with positive affect[schwartz1997neuroanatomical, jatupaiboon2013real]. Notable $\beta$ band enhancement occurred across 6 of electrodes ($p<0.05$, $d=0.29$ to $0.58$), concentrated in Left temporal and frontal regions.
Surprise: Surprise emotion elicited moderate $\alpha$ desynchronization across 21 of electrodes ($p<0.05$, d=0.67d=-0.67 to 0.31-0.31), primarily in parietal and occipital regions, indicating selective posterior cortical activation. Significant $\beta$ band desynchronization was observed across 21 of electrodes ($p<0.05$, d=0.52d=-0.52 to 0.31-0.31), concentrated in temporal and parietal areas, reflecting modulation of sensorimotor and attention networks. Notable $\gamma$ band desynchronization occurred across 16 of electrodes ($p<0.05$, d=0.48d=-0.48 to 0.31-0.31), with clusters in temporal and parietal regions, suggesting high-frequency neural suppression.
Disgust: Disgust emotion elicited moderate $\alpha$ desynchronization across 15 of electrodes ($p<0.05$, d=0.51d=-0.51 to 0.34-0.34), distributed across frontal, central, and parietal regions, indicating widespread cortical engagement. Significant $\theta$ band desynchronization was observed across 6 of electrodes ($p<0.05$, d=0.44d=-0.44 to 0.33-0.33), primarily concentrated in frontal areas, reflecting cognitive control network activation. Minimal $\beta$ band desynchronization was evident in only 2 electrodes ($p<0.05$, d=0.60d=-0.60 to 0.56-0.56), confined to parietal regions.
Anger: Anger emotion elicited extensive $\alpha$ desynchronization across 27 of electrodes ($p<0.05$, d=0.68d=-0.68 to 0.31-0.31), distributed across frontal, temporal, central, and parietal regions, indicating widespread cortical engagement. Significant $\beta$ band desynchronization was observed across 9 of electrodes ($p<0.05$, d=0.49d=-0.49 to 0.30-0.30), primarily in temporal, central, and parietal areas, suggesting sensorimotor network activation associated with action preparation.
Fear: Fear emotion elicited profound $\alpha$ desynchronization across 54 of electrodes ($p<0.05$, d=0.74d=-0.74 to 0.18-0.18), distributed in frontal, temporal, central, and parietal regions, indicating widespread cortical arousal and heightened vigilancet[chien2017oscillatory, bacigalupo2022alpha]. Significant $\beta$ desynchronization was observed across 18 of electrodes ($p<0.05$, d=0.73d=-0.73 to $0.28$), distributed in frontal, central, and parietal regions, suggesting sensorimotor network engagemen[roxburgh2023neural]. Moderate $\theta$ desynchronization occurred across 12 of electrodes ($p<0.05$, d=0.47d=-0.47 to $0.06$), concentrated in central and parietal regions.
Sadness: Sadness emotion elicited extensive $\alpha$ desynchronization across 48 of electrodes ($p<0.05$, d=0.63d=-0.63 to 0.33-0.33), distributed in frontal, temporal, parietal, and occipital regions, indicating widespread cortical engagement. Significant $\beta$ desynchronization was observed across 23 of electrodes ($p<0.05$, d=0.57d=-0.57 to 0.31-0.31), primarily in frontal, temporal, and parietal regions, suggesting sensorimotor network activation. Notable $\gamma$ band desynchronization occurred across 10 of electrodes ($p<0.05$, d=0.46d=-0.46 to 0.31-0.31), concentrated in parietal and occipital regions, reflecting high-frequency neural suppression.
In summary, our research systematically delineates unique neurophysiological signatures for basic emotions using precisely annotated event markers. Widespread $\alpha$ desynchronization served as a common index of cortical arousal, with varying intensity across emotions. Among them, representative examples are Fear demonstrated the most extensive $\alpha$ suppression (54 electrodes) with concurrent $\beta$ and $\theta$ desynchronization; Happiness uniquely featured moderate $\alpha$ suppression (40 electrodes) alongside prominent $\gamma$ and $\beta$ enhancement, reflecting positive affect and reward processing; and Disgust exhibited the most constrained response (15 electrodes) with selective $\alpha$ and $\theta$ desynchronization in frontal regions, suggesting targeted cognitive control activation. These differential oscillatory patterns validate the effectiveness of precise annotation paradigms in capturing genuine neurophysiological correlates of emotion.
We also performed this cluster-based permutation analysis using adjacent window durations 3s and 5s. As shown in the Appendix Fig. 1 and Fig. 2, alternative window durations resulted in a weakening of the key neural signatures. Most notably, in the 5s window ($t_{event}$ - 2s, $t_{event}$ + 3s]), the significant neural patterns for ’Surprise’ and ’Disgust’ disappeared entirely, while the patterns for ’Fear’ $\alpha$ suppression and ’Happiness’ $\gamma$ enhancement were markedly diminished. This finding—that the physiological signal is tightly time-locked to our $t_{\text{event}}$ anchor and degrades as the window shifts or lengthens—demonstrates the precision of our annotation paradigm and validates our 4s window as the most effective choice for this analysis.